OpenStudy (anonymous):

Find the distance between the points given. (-10, 3) and (-10, 12) √(17) 9 15

OpenStudy (anonymous):

\[\sqrt{(-10+10)^2+(12-3)^2}\] \[\sqrt{(0)^2+(9)^2}\] d = 9